What is the molarity of 200.0 mL of solution in which 205.8 g of sodium bromide...

  1. What is the molarity of 200.0 mL of solution in which 205.8 g of sodium bromide is dissolved
Answer #1

The molarity of a solution can be calculated from the formula of molarity which is given as:

  • M = (Wsolute * 1000) / (Vsolution * MWsolute) ......(I)

Here,

M = Molarity of solution

Wsolute = weight of solute

MWsolute = Molecular weight of solute

Vsolution = Volume of solution

Given data:

Solute = NaBr

Wsolute = 205.8 g

Vsolution = 200 mL

The molecular weight of NaBr is 102.89 g/mol.

=> MWsolute (NaBr) = 102.89 g/mol

Substituting the given values in equation (I), we get

M = (Wsolute * 1000) / (Vsolution * MWsolute)

=> M = 205.8 * 1000 / 200 * 102.89

=> M = 205800 / 20578 = 10.0 M

=> Molarity of solution = 10.0 mol/L

Therefore, the molarity of solution is 10.0 M.